# DualXDA: Towards Sparse, Efficient and Explainable Data Attribution in Large AI Models

### v2

@@ Abstract
Contemporary deep learning models achieve remarkable performance over a wide range of domains, yet their decision-making processes often remain opaque. In response, the field of eXplainable Artificial Intelligence (XAI) has grown significantly over the last decade, primarily focusing on feature attribution methods to shed light on which input features drive model predictions. Complementing this perspective, Data Attribution (DA) has emerged as a promising paradigm that shifts the focus from features to data provenance. With the insights gained on the level of (training) data points, DA provides transparency about the model and individual predictions, e.g. for model debugging, identifying data-related causes of suboptimal performance, such as mislabelled instances, dataset distillation or knowledge discovery purposes. However, existing DA approaches suffer from prohibitively high computational costs and memory demands when applied to large-scale or even medium-scale datasets and models, forcing practitioners to resort to approximations that may fail to capture the true inference process of the underlying model. Additionally, current attribution methods exhibit low sparsity, resulting in non-negligible attribution scores across a high number of training examples, hindering the discovery of decisive patterns in the data. In this work, we introduce DualXDA, a framework for sparse, efficient and explainable DA, comprised of two interlinked approaches for Dual Data Attribution (DualDA) and eXplainable Data Attribution (XDA): With DualDA, we propose a novel approach for efficient and effective DA, leveraging Support Vector Machine theory to provide fast and naturally sparse data attributions for AI predictions. In extensive quantitative analyses, we demonstrate that DualDA achieves high attribution quality, excels at solving a series of evaluated downstream tasks, while at the same time improving explanation time by a factor of up to compared to the original Influence Functions method, and up to compared to the method’s most efficient approximation from literature to date. We further introduce XDA, a method for enhancing Data Attribution with capabilities from feature attribution methods to explain why training samples are relevant for the prediction of a test sample in terms of impactful features, which we showcase and verify qualitatively in detail. Taken together, our contributions in DualXDA ultimately point towards a future of eXplainable AI applied at unprecedented scale, enabling transparent, efficient and novel analysis of even the largest neural architectures – such as LLMs – and fostering a new generation of interpretable and accountable AI systems. The implementation of our methods, as well as the full experimental protocol, is available on github111 https://github.com/gumityolcu/DualXDA.
Keywords Explainable AI, Data Attribution, Feature Attribution, Interpretability
@@ 1 Introduction
Despite the remarkable achievements of deep learning approaches, these methods remain black-boxes due to the opacity of their information processing procedures. Explainable AI (XAI) (Samek et al., 2019) has emerged to address the need to elucidate the inference processes of these models. Specifically, global XAI aims to provide insights into overall model characteristics, while local XAI explains outputs for individual test samples. Early approaches have been centered on highlighting relevant input features, with feature attribution methods (e.g. (Bach et al., 2015; Ribeiro et al., 2016)) assigning an attribution score to each input feature, which provides a ranking or a heatmap that indicates which part of the input is more or most influential in determining the model output. Later approaches provide different explanation paradigms, such as concept-based (Kim et al., 2018; Achtibat et al., 2023) and counterfactual explanations (Guidotti, 2024), or offer a comprehensive view on the predictions, model and data (Dreyer et al., 2025).
@@
Feature-centered approaches to XAI put a strong emphasis on model inputs, parameters and representations but ignore the training data that the model has observed during optimization. A relatively recent lens of interpretability, Data Attribution (DA) attributes the model outputs to samples of the training dataset. Complementarily to feature attribution, DA produces attribution scores for each (training) datapoint, allowing one to rank that data with respect to their effect in determining the model parameters (global DA) or specific outputs (local DA).
@@
Data quality is integral in order to obtain robust model performance: Biased sampling (Lapuschkin et al., 2019), poor data quality management (Lin et al., 2022) or adversarial attacks (Shafahi et al., 2018; Zhu et al., 2019), can drive the model to suboptimal solutions. Therefore, in addition to understanding and debugging the decision processes of the model, DA methods have found applications in identifying and solving data-related problems, such as detecting mislabeled samples (Koh & Liang, 2017; Yeh et al., 2018; Pruthi et al., 2020), designing (Fang et al., 2020) and counteracting backdoor attacks (Hammoudeh & Lowd, 2022). Furthermore, these attribution methods have demonstrated effectiveness in dataset distillation (Liu et al., 2021; Naujoks et al., 2024), i.e. finding small subsets of training data which can be used to optimize on without having to sacrifice model performance. This is an important step towards the creation of efficient training loops for deep learning models, which usually require large amounts of data and energy for large scale training in order to be effective predictors (Thompson et al., 2020).
@@
In practice, the utility of most DA methods is limited by two key factors: First, state-of-the-art methods depend on computations that involve the estimation of inverse Hessian vector products, which is computationally expensive due to high runtime and memory demands. While recent work has improved the efficiency of several approaches to DA with the intent to make them feasible for applications on larger language models such as BART or BERT (Guo et al., 2021; Ladhak et al., 2022), these methods remain challenging to scale and are often impractical under strict computational constraints, limiting their accessibility in real-world applications. Some methods can be adjusted to reduce computational load by resorting to lower quality estimations of attributions. Still, the application of these methods in time-critical scenarios, where attributions are required in the order of seconds or less following inference, is exceedingly difficult due to the tightly imposed temporal constraints contrasting their computational complexity.
@@
Secondly, many existing DA methods produce dense attribution scores, i.e., they assign non-negligible importance to a large portion of the training dataset. In contrast, sparse attributions, where only a small subset of candidate components receive substantial credit, can lead to explanations of lower complexity, which have been shown to enhance human interpretability (Colin et al., 2022). The connection between sparsity and interpretability has been well-established in feature attribution literature, where sparse explanations mitigate cognitive load and improve the users’ ability to understand model behavior (Chalasani et al., 2020; Warnecke et al., 2020; Bhatt et al., 2021). This desideratum receives empirical support from Ilyas et al. (Ilyas et al., 2022), who demonstrate that ground-truth attributions are inherently sparse, with only a small subset of training samples exerting influence on the model output predicted on individual test points. To enforce the effect of attribution sparsity, an additional (artificial) sparsification step is sometimes applied to non-sparse methods (Park et al., 2023). This process discards a large portion of the attributions, which in turn, may potentially result in attributions that do not reflect the actual prediction patterns of the model being explained. Regarding explanations of black-box models, the property of closely representing the inference of a model it aims to explain is referred to as “faithfulness” in XAI literature (Hedström et al., 2023). Therefore, sparse attributions that are also faithful to the model are highly preferred in DA.
@@
To address both problems, we turn to established machine learning algorithms, where the influence of individual training data points has already been studied in detail. Concretely, we build upon the concept of Support Vector Machines (SVMs) (Cortes & Vapnik, 1995). SVMs perform linear classification by optimizing weight vectors in terms of a linear combination of a small subset of the training data. This allows the derivation of an alternative formulation of the prediction function as a sum of independent contributions relating to each training sample, by solving the associated dual problem (Crammer & Singer, 2001). Kernel SVMs, in turn, learn linear classifiers on nonlinear representations of training samples, which are then used to perform inference on test data.
Therefore, given a neural network predictor and assuming the penultimate layer thereof as a mapping into a reproducing kernel Hilbert space, we build a SVM-based surrogate model replacing the final layer of the network, yielding an interpretable and faithful surrogate for the whole model.
@@
We also address another limitation of DA methods: the attribution values themselves are opaque, providing no insight into why certain training datapoints are deemed important for a given prediction. To this end, we combine data attribution with feature attribution to identify relevant features that explain the attribution values in both training and test input domains simultaneously.
@@
We introduce DualDA: a novel DA method which employs multiclass kernel SVMs (Crammer & Singer, 2001) as surrogates for deep learning models. Based on their dual problem, we derive sparse and high-quality explanations, both locally and globally, in a time- and memory-efficient manner.
We evaluate DualDA in terms of quality and sparsity of explanations, and runtime as well as memory efficiency. In total, we compare nine different methods on seven quantitative metrics across three different datasets and models, in the most extensive quantitative evaluation of prominent DA methods so far. We find that DualDA performs at state-of-the-art while providing an up to 4,100,000 speed-up compared to the least efficient competing method and up to 11,000 compared with the SoTA implementations of Influence Functions. By making data attribution tractable at scale, our approach opens up a broad range of new practical applications – from large model auditing to interactive dataset diagnostics – that were previously infeasible due to computational constraints.
We prove that DualDA attribution scores agree within the multiclass kernelized SVM setting, on almost all datapoints, with attributions produced by both the gradient-based Influence Functions framework (Koh & Liang, 2017) (with minor modification) and the kernel surrogate-based Representer Points framework (Yeh et al., 2018).
We also introduce XDA: a method that leverages the popular Layer-wise Relevance Propagation (LRP) (Bach et al., 2015; Montavon et al., 2019; Samek et al., 2021; Achtibat et al., 2024) – a state-of-the-art feature attribution (FA) method – and eXplains attribution values assigned to samples by mapping them back through the model to the (input) feature space. This yields insights as to why a certain training datapoint is important by identifying relevant features – such as pixels or tokens – both in the training and test input domains simultaneously. To our knowledge, XDA is the first method to enable a seamless integration of feature attribution and data attribution, offering joint explanations that connect model behaviour across feature and data provenance dimensions.
@@
In summary, our contributions address the core scalability bottleneck of data attribution by introducing DualDA, a highly efficient and sparse method that enables practical deployment of DA across large models and datasets. Building on this foundation, we present XDA as a bridge to feature attribution, enabling – for the first time – joint, interpretable explanations that connect influential training samples to meaningful input features. Together, these components form the DualXDA framework, providing novel opportunities for deeper, more actionable analyses into the inner workings of modern neural networks.
@@ 2.1 What is Data Attribution?
Global DA methods assign each pair of training sample and class a real-valued attribution which indicates the relevance or influence of the training datapoint on the fit of the model . Local DA methods produce a function for each class which accepts as input a test sample and the index for a training point. The value indicates the relevance or influence of the training datapoint on the model output .
DA has historically been framed as the problem of approximating the ground-truth effects of model retraining under altered training data compositions (Koh & Liang, 2017; Park et al., 2023). Recent work (Wei et al., 2024) advocates for a different standpoint, aiming to estimate the changes induced by fine-tuning an already trained model with an altered training set instead. Other authors do not claim for their attribution values to approximate a ground truth, instead they are interested in the utility of their attribution values in fulfilling certain downstream tasks. In all of these formulations, the resulting attributions are interpreted such that positive attribution means the training sample contains evidence for the model decision while negative attribution is indicative of evidence against the decision being explained.
@@ 2.2 Data Attribution Methods
Existing DA methods can be broadly categorized into four different groups, in terms of how they produce attribution scores:
@@
Similarity-based approaches compute the similarity between the training and test samples, building on the intuition that the model inferences should be informed of training samples that produce similar latent representations (Pezeshkpour et al., 2021), or through other measures of similarity (Singla et al., 2023). However, these attributions are class-unspecific as the measured similarity is independent of the predicted class.
@@
Gradient-based approaches estimate the effect of retraining on a new training set where a sample has been removed from the training corpus. In this framework, a positive attribution indicates a decrease in loss, which indicates more confidence for the associated class. To this end, these methods employ Taylor approximations, which involve gradients of the training and test sample. The earliest method, Influence Functions (Koh & Liang, 2017), calculates a first-order approximation using the parameters of the final trained model. As this involves calculations of the inverse Hessian, which are infeasible to compute in limited time for even modestly-sized models, approximations using the LiSSA algorithm (Agarwal et al., 2017), the Arnoldi method (Schioppa et al., 2022) or the (Eigenvalue-corrected) Kronecker-Factored Approximate Curvature method (EK-FAC) (Grosse et al., 2023) are commonly used. Furthermore, the method TRAK (Park et al., 2023) applied to a single model can be seen as a version of Influence Functions with a general Gauss-Newton approximation of the Hessian. To include the behaviour of the model over the entire training process, the method TracIn (Pruthi et al., 2020) collects the dot product between the gradients at different epochs and accumulates them in a weighted manner in relation to the corresponding step sizes. GradDot and GradCos (Charpiat et al., 2019) are simplifications of TracIn which only consider the (normalised) dot product at the final epoch.
@@
Surrogate-based approaches propose exchanging parts of the network for similarly behaving surrogates which are inherently interpretable in terms of the importance of training data points for predictions. The Representer Points (Yeh et al., 2018) method retrains only the final fully-connected layer of the model using a modified loss with an added -penalty. This enables the application of the Representer Theorem to decompose the networks output into the sum of kernel evaluations involving the test point and the individual training points.
@@
Retraining-based approaches sample counterfactual models by retraining the model on various altered training sets and then use these samples to estimate the model output for an arbitrary training subset. Consequently, these methods are generally regarded as baselines and are very demanding in terms of computational resources (Park et al., 2023).
Detailed explanations and formulas for similarity-, surrogate- and gradient-based methods can be found in Section D.1.
@@ 3 Dual eXplainable Data Attribution
DualDA is a surrogate-based DA method, which replaces the final linear layer of the model with a linear multiclass SVM. We selected SVMs as surrogates for four reasons: (1) they closely approximate the behavior of the original model, (2) are inherently interpretable in a DA sense in their dual form, (3) yield sparse attributions, (4) and enable fast training and inference in large AI models, as further outlined below.
(1) Prior research shows that gradient descent in various models has an implicit bias towards margin maximizing behaviour (Gunasekar et al., 2018; Soudry et al., 2018; Gidel et al., 2019; Ji & Telgarsky, 2019; Tarzanagh et al., 2023) which is the defining feature of SVMs. In particular, recent studies demonstrate convergence of neural networks to kernelized SVMs under certain conditions in binary and multiclass classification scenarios (Lyu & Li, 2019; Ravi et al., 2024). This provides the theoretical motivation to assume that an SVM can serve as a faithful surrogate for the original final layer of a neural network predictor. We validate this assumption empirically in Section 4.5.
@@
(3) When working with large datasets containing millions of samples, attribution methods that yield similar or indistinguishable attributions across a large number of data points become difficult to interpret effectively (Iwata & Yoshikawa, 2021). SVMs exhibit an implicit bias towards sparsity in their attribution values as, under constrained conditions, only a limited number of support vectors contribute in determining the model’s decision boundary. While this sparsity can enhance interpretability by focusing on the most influential samples, excessive sparsity may hinder the expressiveness of the resulting explanations. To mitigate this, the degree of sparsity can be controlled through the choice of the regularization hyperparameter of the underlying SVM, allowing a balance between conciseness and representational richness in the attributions, as shown in detail in Section F.3.
(4) The training of a linear SVM has a worst-case time complexity between and (Hush et al., 2006), where is the size of the attribution (e.g. training) dataset and is the size of the input space of the SVM, i.e. the size of the penultimate layer. This allows efficient applicability on large datasets and models. Additionally, computing attributions for a single pair of train and test samples requires only an inner product. The computational complexity per local attribution score is therefore only . This enables near-instantaneous attribution calculation for new test samples after caching hidden layer activations of the attribution dataset and SVM weights.
@@
A further extension of our approach including explanations using human understandable concepts to fully unify explanations on the data-level, feature-level and concept-level ( inspired by (Achtibat et al., 2023) ) is given in Appendix H.

### v3

# Sparse, Efficient and Explainable Data Attribution with DualXDA

@@ Abstract
Data Attribution (DA) is an emerging approach in the field of eXplainable Artificial Intelligence (XAI), aiming to identify influential training datapoints which determine model outputs. It seeks to provide transparency about the model and individual predictions, e.g. for model debugging, identifying data-related causes of suboptimal performance. However, existing DA approaches suffer from prohibitively high computational costs and memory demands when applied to even medium-scale datasets and models, forcing practitioners to resort to approximations that may fail to capture the true inference process of the underlying model. Additionally, current attribution methods exhibit low sparsity, resulting in non-negligible attribution scores across a high number of training examples, hindering the discovery of decisive patterns in the data. In this work, we introduce DualXDA, a framework for sparse, efficient and explainable DA, comprised of two interlinked approaches, Dual Data Attribution (DualDA) and eXplainable Data Attribution (XDA): With DualDA, we propose a novel approach for efficient and effective DA, leveraging Support Vector Machine theory to provide fast and naturally sparse data attributions for AI predictions. In extensive quantitative analyses, we demonstrate that DualDA achieves high attribution quality, excels at solving a series of evaluated downstream tasks, while at the same time improving explanation time by a factor of up to compared to the original Influence Functions method, and up to compared to the method’s most efficient approximation from literature to date. We further introduce XDA, a method for enhancing Data Attribution with capabilities from feature attribution methods to explain why training samples are relevant for the prediction of a test sample in terms of impactful features, which we showcase and verify qualitatively in detail. Taken together, our contributions in DualXDA ultimately point towards a future of eXplainable AI applied at unprecedented scale, enabling transparent, efficient and novel analysis of even the largest neural architectures – such as Large Language Models – and fostering a new generation of interpretable and accountable AI systems. The implementation of our methods, as well as the full experimental protocol, is available on github111 https://github.com/gumityolcu/DualXDA.
@@ 1 Introduction
Despite the remarkable achievements of deep learning approaches, these methods remain black-boxes due to the opacity of their information processing procedures. Explainable AI (XAI) (samek2019explainable) has emerged to address the need to elucidate the inference processes of these models. Specifically, global XAI aims to provide insights into overall model characteristics, while local XAI explains outputs for individual test samples. Early approaches have been centered on feature attribution (misc:xai:LRP; misc:LIME), assigning an attribution score to each input feature, which indicates which part of the input is more or most influential in determining the model output. Later approaches provide different explanation paradigms, such as concept-based (kim2018interpretability; achtibat2023attribution) and counterfactual explanations (guidotti2024counterfactual).
@@
Feature-centered approaches put a strong emphasis on model inputs, parameters and representations but ignore the training data that the model has observed during optimization. A relatively recent lens of interpretability, Data Attribution (DA) attributes the model outputs to samples of the training dataset, producing attribution scores for each (training) datapoint. This is a complementary approach to feature attribution, allowing to rank the training data with respect to their effect in determining the model parameters (global DA) or specific outputs (local DA).
@@
Data quality is integral in order to obtain robust model performance: Biased sampling (misc:xai:cleverhans), poor data quality management (lin2022measuring) or adversarial attacks (shafahi2018poison; zhu2019transferable), can drive the model to suboptimal solutions. Therefore, in addition to understanding and debugging the decision processes of the model, DA methods have found applications in identifying and solving data-related problems, such as detecting mislabeled samples (method:koh_influence; method:representer; method:trackin), designing (fang2020influence) and counteracting backdoor attacks (hammoudeh2022identifying). Furthermore, these attribution methods have demonstrated effectiveness in dataset distillation (liu2021influence; naujoks2024pinnfluence), i.e. finding small subsets of training data which can be used to optimize on without having to sacrifice model performance. This is an important step towards the creation of efficient training loops for deep learning models, which usually require large amounts of data and energy for large scale training in order to be effective predictors (thompson2020computational).
@@
In practice, the utility of most DA methods is limited by two key factors: First, state-of-the-art methods depend on the estimation of inverse Hessian vector products, which is computationally expensive. While recent work has improved the efficiency of several approaches to DA with the intent to make them feasible for applications on larger language models such as BART or BERT (guo2020fastif; ladhak2022contrastive), these methods remain challenging to scale and are often impractical under strict computational constraints, limiting their accessibility in real-world applications. Some methods can be adjusted to reduce computational load by resorting to lower quality estimations of attributions. Still, the application of these methods in time-critical scenarios, where attributions are required in the order of seconds or less following inference, is exceedingly difficult due to the tightly imposed temporal constraints contrasting their computational complexity.
@@
Secondly, many existing DA methods produce dense attribution scores, i.e., they assign non-negligible importance to a large portion of the training dataset. In contrast, sparse attributions, where only a small subset of candidate components receive substantial credit, can lead to explanations of lower complexity, which have been shown to enhance human interpretability (colin2022cannot). The connection between sparsity and interpretability has been well-established in feature attribution literature, where sparse explanations mitigate cognitive load and improve the users’ ability to understand model behavior (chalasani2020concise; warnecke2020evaluating; bhatt2020evaluating). This desideratum receives empirical support from Ilyas et al. (ilyas2022datamodels), who demonstrate that ground-truth data attributions are inherently sparse, with only a small subset of training samples exerting influence on the model output predicted on individual test points. To enforce the effect of attribution sparsity, an additional (artificial) sparsification step is sometimes applied to non-sparse methods (misc:TRAK). This process discards a large portion of the attributions, which in turn, may potentially result in attributions that do not reflect the actual prediction patterns of the model being explained. Regarding explanations of black-box models, the property of closely representing the inference of a model it aims to explain is referred to as “faithfulness” in XAI literature (hedstrom2023quantus). Therefore, sparse attributions that are also faithful to the model are highly preferred in DA.
@@
To address both problems, we turn to established machine learning algorithms, where the influence of individual training data points has already been studied in detail. Concretely, we build upon the concept of Support Vector Machines (SVMs) (cortes1995support). SVMs perform linear classification by optimizing weight vectors in terms of a linear combination of a small subset of the training data. This allows the derivation of an alternative formulation of the prediction function as a sum of independent contributions relating to each training sample, by solving the associated dual problem (svm:CS). Kernel SVMs, in turn, learn linear classifiers on nonlinear representations of training samples, which are then used to perform inference on test data. Therefore, given a neural network predictor and assuming the penultimate layer thereof as a mapping into a reproducing kernel Hilbert space, we build a SVM-based surrogate model replacing the final layer of the network, yielding an interpretable and faithful surrogate for the whole model.
@@
Finally, we address another limitation of DA methods: the attribution values themselves are opaque, providing no insight into why certain training datapoints are deemed important for a given prediction. To this end, we combine data attribution with feature attribution to identify relevant features that explain the attribution values in both training and test input domains simultaneously.
@@
We introduce DualDA: a novel DA method which employs multiclass kernel SVMs (svm:CS) as surrogates for deep learning models. Based on their dual problem, we derive sparse and high-quality explanations, both locally and globally, in a time- and memory-efficient manner.
We evaluate DualDA in terms of quality and sparsity of explanations, and runtime as well as memory efficiency. In total, we compare nine different methods on seven quantitative metrics across three different datasets and models, in the most extensive quantitative evaluation of prominent DA methods so far. We find that DualDA performs at state-of-the-art while providing an up to 4,100,000 speed-up compared to the least efficient competing method and up to 11,000 compared with the SoTA implementations of Influence Functions. These efficiency gains meaningfully expand the practical usability of data attribution, including scenarios where existing methods are too computationally demanding.
We prove that DualDA attribution scores agree within the multiclass kernelized SVM setting, on almost all datapoints, with attributions produced by both the gradient-based Influence Functions framework (method:koh_influence) (with minor modification) and the kernel surrogate-based Representer Points framework (method:representer).
We also introduce XDA: a method that leverages the popular Layer-wise Relevance Propagation (LRP) (misc:xai:LRP; montavon2019layer; samek2021explaining; achtibat2024attnlrp) – a state-of-the-art feature attribution (FA) method – and eXplains attribution values assigned to samples by mapping them back through the model to the (input) feature space. This yields insights as to why a certain training datapoint is important by identifying relevant features – such as pixels or tokens – both in the training and test input domains simultaneously. To our knowledge, XDA is the first method to enable a seamless integration of feature attribution and data attribution, offering joint explanations that connect model behaviour across feature and data provenance dimensions.
@@
In summary, we propose DualDA as an efficient, sparse and effective alternative to existing Data Attribution methods. Building on this foundation, we present XDA as a bridge to feature attribution, enabling – for the first time – joint, interpretable explanations that connect influential training samples to meaningful input features. Together, these components form the DualXDA framework, providing novel opportunities for deeper, more actionable analyses into the inner workings of modern neural networks.
@@ 2.1 What is Data Attribution?
Global DA methods assign each pair of training sample and class a real-valued attribution which indicates the relevance or influence of the training datapoint on the fit of the model . Local DA methods produce a function for each class . The inputs to this function are a test sample and the index for a training point. The value indicates the relevance or influence of the training datapoint on the model output .
DA has historically been framed as the problem of approximating the ground-truth effects of model retraining under altered training data compositions (method:koh_influence; misc:TRAK). Recent work (wei2024final) advocates for a different standpoint, aiming to estimate the changes induced by fine-tuning an already trained model with an altered training set instead. Other authors do not claim for their attribution values to approximate a ground truth, instead they are interested in the utility of their attribution values in fulfilling certain downstream tasks. In all of these formulations, the resulting attributions are interpreted such that positive attribution means the training sample contains evidence for the model decision while negative attribution is indicative of evidence against the decision being explained.
@@ 2.2 Data Attribution Methods
Existing DA methods can be broadly categorized into four different groups, in terms of how they produce attribution scores, as listed below. Detailed explanations and formulas for all methods can be found in Section D.1.
@@
Similarity-based approaches compute the similarity between the training and test samples, building on the intuition that the model inferences should be informed of training samples that produce similar latent representations (pezeshkpour2021empirical), or through other measures of similarity (singla2023simple). However, these attributions are class-unspecific as the measured similarity is independent of the predicted class.
@@
Gradient-based approaches estimate the effect of retraining on a new training set where a sample has been removed from the training corpus. In this framework, a positive attribution indicates a decrease in loss, which indicates more confidence for the associated class. To this end, these methods employ Taylor approximations, which involve gradients of the training and test sample. The earliest method, Influence Functions (method:koh_influence), calculates a first-order approximation using the parameters of the final trained model. As this involves calculations of the inverse Hessian, which are infeasible to compute in limited time for even modestly-sized models, approximations using the LiSSA algorithm (misc:influence_agarwal), the Arnoldi method (schioppa2022scaling) or the (Eigenvalue-corrected) Kronecker-Factored Approximate Curvature method (EK-FAC) (grosse2023studying) are commonly used. Furthermore, the method TRAK (misc:TRAK) applied to a single model can be seen as a version of Influence Functions with a general Gauss-Newton approximation of the Hessian. To include the behaviour of the model over the entire training process, the method TracIn (method:trackin) collects the dot product between the gradients at different epochs and accumulates them in a weighted manner in relation to the corresponding step sizes. GradDot and GradCos (method:perspective) are simplifications of TracIn which only consider the (normalised) dot product at the final epoch.
@@
Surrogate-based approaches propose exchanging parts of the network for similarly behaving surrogates which are inherently interpretable in terms of the importance of training data points for predictions. The Representer Points (method:representer) method retrains only the final fully-connected layer of the model using a modified loss with an added -penalty. This enables the application of the Representer Theorem to decompose the networks output into the sum of kernel evaluations involving the test point and the individual training points.
@@
Retraining-based approaches sample counterfactual models by retraining the model on various altered training sets and then use these samples to estimate the model output for an arbitrary training subset. Consequently, these methods are generally regarded as baselines and are very demanding in terms of computational resources (misc:TRAK).
@@ 3 Dual eXplainable Data Attribution
DualDA is a surrogate-based DA method, which replaces the final linear layer of the model with a linear multiclass SVM. We selected SVMs as surrogates for four reasons: (1) they closely approximate the behavior of the original model, (2) are inherently interpretable in a DA sense in their dual form, (3) yield sparse attributions, (4) and enable fast training and inference, as further outlined below.
(1) Prior research shows that gradient descent in various models has an implicit bias towards margin maximizing behaviour (bias:convolutional; bias:logistic; bias:linear2; bias:linear; bias:transformers) which is the defining feature of SVMs. In particular, recent studies demonstrate convergence of neural networks to kernelized SVMs under certain conditions in binary and multiclass classification scenarios (bias:homogeneous; ravi2024implicit). This provides the theoretical motivation to assume that an SVM can serve as a faithful surrogate for the original final layer of a neural network predictor. We validate this assumption empirically in Section 4.5.
@@
(3) When working with large datasets containing millions of samples, attribution methods that yield similar or indistinguishable attributions across a large number of data points become difficult to interpret effectively (misc:sparse_tda). SVMs exhibit an implicit bias towards sparsity in their attribution values as, under constrained conditions, only a limited number of support vectors contribute in determining the model’s decision boundary. While this sparsity can enhance interpretability by focusing on the most influential samples, excessive sparsity may hinder the expressiveness of the resulting explanations. To mitigate this, the degree of sparsity can be controlled through the choice of the regularization hyperparameter of the underlying SVM, allowing a balance between conciseness and representational richness in the attributions, as shown in detail in Section F.3.
(4) The training of a linear SVM has a worst-case time complexity between and (hush2006qp), where is the size of the attribution (e.g. training) dataset and is the size of the input space of the SVM, i.e. the size of the penultimate layer. This allows efficient applicability on large datasets and models. Additionally, computing attributions for a single pair of train and test samples requires only an inner product. The computational complexity per local attribution score is therefore only . This enables near-instantaneous attribution calculation for new test samples after caching hidden layer activations of the attribution dataset and SVM weights.
@@
A further extension of our approach including explanations using human understandable concepts to fully unify explanations on the data-level, feature-level and concept-level ( inspired by (achtibat2023attribution) ) is given in Appendix H.